I need to figure out how to use point slope form to find the equation of the line with given properties and write the equation in general form. m= -7 passing through(3,0) = 21. thank you
Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

I need to figure out how to use point slope form to find the equation of the line with given properties and write the equation in general form. m= -7 passing through(3,0) = 21.
I don't know what the 21 means ??
--------------
Do it like this one.
------------------------
The slope, m = -3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note.

Answer by richwmiller(17219)

You almost had it but this (3,0)=21 means nothing.
y=mx+b
0=3*-7+b
21=b
y=-7x+21
Perhaps you meant that the y intercept (b) =21
y=-7x+21 is in slope intercept form
Standard form is ax+by+c=0
subtract y from both sides
0=-7x-y+21
-7x-y+21=0